Identify the graph of each of the following showing the centre, foci, length of the latus rectum, the eccentricity and asymptotes where applicable

I. 25x² + 4y² = 100

II. 25x² + 16y² = 400


1
Expert's answer
2022-03-20T06:42:03-0400

i) 25x² +4y²=100 can be written as x²/4 + y²/25= –1

This is of the form x²/a² –y²/b² –1

therefore, a²= 4, b²= 25 and giving us

a= 2, b=5

Lenth of transverse axis: 2b= 2(5)=10

length of conjugate axis: 2a= 2(2)= 4

Eccentricity: √{1+a²/b²)==> √ {1+4/25}

=√29/5

Foci: The co- ordinate of the foci are (0,±be) i.e (0,+√29)

Vertices: The co–ordinate of the vertices are ( 0,± be) ( 0,±5)

Asymptotes y = ± b/e giving ±25/√29

ii) 25x² +16y² =400 can be written as x²/16 + y²/25= –1

This is of the form x²/a² –y²/b² –1

therefore, a²= 16, b²= 25 and giving us

a= 4, b=5

Lenth of transverse axis: 2b= 2(5)=10

length of conjugate axis: 2a= 2(4)= 8

Eccentricity: √{1+a²/b²)==> √ {1+16/25}

=√41/5

Foci: The co- ordinate of the foci are (0,±be) i.e (0,+√41)

Vertices: The co–ordinate of the vertices are ( 0,± be) ( 0,±5)

Asymptotes y = ± b/e giving ± 25/√41